# Meson dynamics from locally exciting a particle-conserving $Z_{2}$ lattice gauge theory

### v1

@@ Abstract
Quantum simulation of lattice gauge theories is an important avenue to gain insights into both particle physics phenomena and constrained quantum many-body dynamics. There is a growing interest in probing analogs of high energy collision phenomena in lattice gauge theories that can be implemented on current quantum simulators. Motivated by this, we characterize the confined mesons that originate from a local high energy excitation in a particle-conserving 1D $Z_{2}$ lattice gauge theory. We focus on a simple, experimentally accessible setting that does not require preparation of colliding wavepackets and isolates the effects of gauge field confinement strength and initial state energy on the nature of propagating excitations. We find that the dynamics is characterized by the propagation of a superposition of differently sized mesons. The linear confinement leads to meson size oscillations in time. The average meson size and oscillation frequency are controlled by the strength of the gauge field confinement. At a constant confinement field, the average meson length is controlled by the initial excitation’s energy. Higher energies produce longer mesons and their effective mass depends strongly on their size: longer mesons propagate more slowly out of the central excitation. Mesons of different sizes get spatially filtered with time due to different speeds. We show that this phenomenology is a consequence of linear confinement and remains valid in both the strong and weak confinement limit. We present simple explanations of these phenomena supported by exact numerics.
@@ I Introduction
Quantum computers and quantum simulators can bridge this gap. By implementing these gauge theories on a discrete lattice, quantum computers can exactly simulate real-time quantum dynamics of high-energy phenomena in a table-top experiment. In recent years, there has been considerable progress in quantum simulation experiments of some simple models such as $Z_{2}$ and $U(1)$ lattice gauge theories [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
@@
The study of dynamical phenomena in lattice gauge theories is attracting considerable attention in recent years. These gauge theories show exotic non-ergodic phenomena such as lack of thermalization due to confinement and Hilbert-space fragmentation [31, 32, 33, 34, 8], quantum many-body scars [35, 36, 37, 38, 39] and Floquet time crystals [40]. Analogs of QCD phenomena such as confinement dynamics [41, 14, 13, 42], pair production through string breaking [43, 44, 17, 45, 46, 47, 15, 16, 12], false vacuum decay [48], and dynamical hadron formation [49, 50] have been found in theoretical calculations and some quantum simulator experiments.
These successes have opened up an exciting new avenue of probing analogs of high energy particle collision phenomena within these simple lattice gauge theories. Some recent works have studied schemes to prepare moving wavepackets for collisions [51, 52] and numerically analyzed scattering processes [53, 54, 55, 49]. In these studies, the dynamics involves a rich interplay of multiple kinds of bound excitations, resonant processes of particle pair production and string fluctuation dynamics.
@@
We also consider the effect of tuning the initial state energy at a fixed confinement field. The average length of mesons that propagate monotonically increases as a function of initial state energy, instead of the spatial size of the initial excitation. As the meson size increases, its propagation speed is slower, which slows down the spread of particle density away from the excitation region. Due to this differential in meson speed, the mesons separate out spatially according to their size as a function of time. This is a consequence of linear confinement and we find that this feature holds true even for cases where the confining field is weaker than the particle hopping term. We show that for weaker confining fields, mesons longer than a certain size experience a slow down.
@@ II Model
Figure 1: (a) Dynamics of mesons: Mesons propagate out of the localized energy excitation in both directions in a 1D chain governed by a $Z_{2}$ lattice gauge theory. The state is described by a superposition of different meson sizes. The confinement field and the initial excitation energy determine the meson size distribution. The propagation speed varies with the meson size. (b) A basis state of the 1D $Z_{2}$ lattice gauge theory model. Open and filled circles represent absence and presence of a boson on sites labeled by index $i$. The red and blue arrows are link gauge spins representing $\sigma^{z}=\pm 1$ respectively. The quantity, $G_{i}=\sigma^{z}_{i-1,i}(-1)^{n_{i}}\sigma^{z}_{i,i+1}=1$ is conserved on each site. The red ellipse denotes a meson where the two bosons are next to each other, called an $r=1$ meson.
@@
There is a local conservation law at each site since $[G_{i},H]=0$. The Hamiltonian eigenstates can thus be chosen to also be the eigenstates of the operator $G_{i}$ at each site. The eigenvalues of $G_{i}$ are ±$\pm 1$. Different choices of the $G_{i}$ eigenvalues correspond to independent and physically distinct sectors of the Hamiltonian’s Hilbert space [24]. Throughout this work, we only consider states $|\psi\rangle$ where $G_{i}|\psi\rangle=+1|\psi\rangle$ for all sites $i$. This constraint implies that $\sigma^{z}_{i-1,i}(-1)^{n_{i}}\sigma^{z}_{i,i+1}|\psi\rangle=|\psi\rangle$. We can rewrite it as, $\sigma^{z}_{i-1,i}\sigma^{z}_{i,i+1}|\psi\rangle=e^{i\pi/2(\sigma^{z}_{i-1,i}-\sigma^{z}_{i,i+1})}=(-1)^{n_{i}}|\psi\rangle$. This can be interpreted as a Gauss law where the local matter density determines the electric field configuration around it.
@@ III The two-particle sector
[…]
@@
$H=-J\sum_{i_{1},i_{2}}(|i_{1}+1,i_{2}\rangle\langle i_{1},i_{2}|+|i_{1},i_{2}+1\rangle\langle i_{1},i_{2}|+\text{h.c.})\\ +2h\sum_{i_{1},i_{2}}|i_{1}-i_{2}||i_{1},i_{2}\rangle\langle i_{1},i_{2}|.$ (2)
We use the center-of-mass $(c=(i_{1}+i_{2})/2)$ and relative $(r=i_{1}-i_{2})$ coordinates to further simplify the Hamiltonian. The Hamiltonian in Eq. 2 becomes
$H=-J\sum_{c}\sum_{r=1}^{\infty}(|c+1/2\rangle\langle c|+\text{h.c.})(|r+1\rangle\langle r|+\text{h.c.})\\ +2h\sum_{r=1}^{\infty}r|r\rangle\langle r|$ (3)
Here the center of mass coordinate $c$ can take half-integer and integer values while the relative coordinate $r$ only takes integer values. The relative coordinate has a linear potential term in Eq. 3. In the ground state, at any finite $h/J$, two particles are confined by this potential into mesons [56]. The center-of-mass coordinate is free and thus the mesons can delocalize. In this work, we study the two-particle sector dynamics in terms of the center-of-mass and relative coordinates. The different relative coordinate values correspond to different types of possible meson excitations. The smallest possible meson is the $r=1$ meson where the two bosons are right next to each other. In later sections, we calculate how the size of the propagating mesons depends on field term $h/J$ and the energy of the initial local excitation. We will also determine how meson size affects its propagation speed.
@@ III.1 Diagonalizing the Hamiltonian
We can diagonalize the Hamiltonian in Eq. 3 by considering the eigenstates to be of the form, $|\psi_{n,k}\rangle=\Phi_{k}\otimes\Gamma_{n}=(\sum_{c}\phi^{k}_{c}|c\rangle)\otimes(\sum_{r}\gamma^{n}_{r}|r\rangle)$ that satisfy the equation, $H|\psi_{n,k}\rangle=E_{n,k}|\psi_{n,k}\rangle$. Note that the center-of-mass states only have a hopping term in the Hamiltonian. Thus plane wave states of the form, $\phi^{k}_{c}=e^{ikc}/\sqrt{2L-1}$ are eigenstates for the center-of-mass degree of freedom. Here $L$ is the total number of sites in the chain. Due to the half integer steps of the coordinate $c$, the momentum $k$ ranges from $-2\pi$ to $2\pi$ in steps of $2\pi/(2L-1)$. We can find the function $\gamma^{n}_{r}$ by substituting $\phi^{k}_{c}$ into the eigenvalue equation of the Hamiltonian in Eq. 3 and projecting it onto $|c\rangle\otimes|r\rangle$. We then get
@@
Note that the relative coordinate is restricted to $r\geq 1$. Without such a restriction, this describes a Wannier-Stark ladder and Eq. 4 admits Bessel function solutions where $\gamma^{n}_{r}=J_{r-n}(2J\cos(k/2)/h)$ [57]. The energy eigenvalue in that case is given by, $E_{n,k}=2hn$ where $n$ is integer quantized. The energy is independent of the center-of-mass momentum leading to a large degeneracy. However in our case, $r\leq 0$ is not allowed and thus we must enforce the boundary condition that $\gamma^{n}_{r=0}=0$. This is akin to adding an infinite potential wall at $r=0$. As a result, the eigenstates are no longer Bessel functions. The exact eigenstates are given by Lommel polynomials and the eigenenergies are obtained from the zeros of the Lommel polynomial [58].
Even with this boundary condition, the eigensolutions are analytically rather simple in the large-$h/J$ and small-$h/J$ limit. These limits essentially capture a complete picture of the physics. We now derive the solution in these two simple limits.
@@ III.1.1 Large-$h/J$ limit
When $h/J\gg 1$, the Bessel function solutions seen earlier approach the exact eigenstates. At large $h/J$, the asymptotic form of the solutions is, $J_{r-n}\sim(2J\cos(k/2)/h)^{|r-n|}$. The amplitude at $r=0$ is $\sim(2J\cos(k/2)/h)^{-|n|}$. As $h/J\to\infty$, when $n\geq 1$, this amplitude goes to zero, satisfying the boundary condition in the asymptotic limit. The energy eigenvalues are quantized in this limit and given by, $E_{n,k}=2hn$. The $n^{\text{th}}$ eigenstate is sharply peaked around $r=n$ with power law tails ($1/h^{|r-n|}$). Seen another way, the zeroth order $h/J\to\infty$ solution has the $n^{\text{th}}$ eigenstate as a delta function on $r=n$ with the eigenenergy $2hn$ where $n\geq 1$. The hopping term proportional to $J$ in Eq. 4 provides first order perturbation to the wavefunction and gives it a tail such that the amplitude on level $r\neq n$ is $(2J\cos(k/2)/h)^{|r-n|}$.
@@ III.1.2 Small-$h/J$ limit
$-\frac{d^{2}\gamma^{n}_{r}}{dr^{\prime 2}}+(r^{\prime}-E^{\prime}_{n,k})\gamma^{n}_{r}=0$ (6)
@@
At intermediate values of the confining field $h/J$, the exact eigenfunctions are given by Lommel polynomials [58]. We will later see that even for $h/J$ as small as $\sim 1$, the functional form of quantities such as average meson size and size oscillation frequencies can be extracted from the large-$h/J$ limit solutions.
@@ IV Effect of confinement field on meson dynamics
In this section, we characterize the meson dynamics from simple initial states as a function of confinement field strength. We first consider the initial excitation where an $r=1$ meson is localized at the center of a 1D chain of length $L$. This corresponds to the initial configuration shown in Fig. 1(b). In the $|r,c\rangle$ (relative and center-of-mass coordinate) basis, the wavefunction at time $t=0$ is given by $|\psi(0)\rangle=|1,(L+1)/2\rangle$. Under the Hamiltonian in Eq. 3, the system wavefunction evolves in time and can be written as, $|\psi(t)\rangle=\sum_{c,r}g_{c,r}(t)|c,r\rangle$. In our calculations, we fix $L=100$ and use exact diagonalization to calculate the time-evolved wavefunctions.
@@
Figure 2: (i),(ii): Particle density as a function of site index (x-axis) and time, $Jt$ (y-axis) for (i) $h/J=1.1$ and (ii) $h/J=0.3$. With time, the mesons delocalize in light-cones like structures. The outermost cones represent faster moving excitations. The slower moving inner cones show substructures with diamond shaped particle tracks denoting temporal oscillation of meson size. (iii),(iv): Snapshot of occupation probability as a function of meson size ($r$) and its center-of-mass position ($c$) at (iii) $Jt=50$ for $h=1.1J$ and (iv) $Jt=30$ for $h=0.3J$. In both cases, smaller sized mesons delocalize the farthest, suggesting larger mesons propagate slower. The different speeds lead to a spatial filtering of meson sizes as larger mesons have more probability of being detected closer to the center during time evolution.
@@
We show the dynamics of propagating excitations in Fig. 2 at two values of the confining field, $h=1.1J$ and $h=0.3J$. In Fig. 2(i),(ii), particles delocalize through light-cone like structures emanating from the central excitation region at time $t=0$. The outermost cones in both cases represent fast moving excitations. The inner cone has a slower spread and has substructures that show diamond shaped particle tracks that also slowly spread out from the center. These are more pronounced and wider for smaller $h=0.3J$. These tracks show similarities to the ones found after wavepacket collisions in Ref. [55, 51] where they were interpreted as the two particles coming together and colliding multiple times. We will show in later sections that we can interpret them as meson size oscillations which gradually spread out from the center as these mesons delocalize. In Fig. 2(iii),(iv), we can clearly see that at a given time, smaller mesons occupy the largest center-of-mass positions away from the central excitation region. This suggests that meson size is inversely related to propagation speed. Due to different speed, mesons of different sizes get spatially filtered in time evolution. We will show that this is a consequence of the linear confinement of mesons and persists for both $h>J$ and $h<J$ regimes.
@@
We use the system wavefunction at long times to characterize the propagating excitations. In particular, we calculate the average meson size, $r_{\text{avg}}(t)=\sum_{r}r|\langle r|\psi(t)\rangle|^{2}$ as a function of time. The meson excitations propagate equally in both the directions away from the center. We determine the speed at which the mesons propagate by calculating the temporal slope of the center-of-mass coordinate’s standard deviation, $c_{s}$, defined as
@@
At time $t=0$, the average center-of-mass coordinate is at $(L+1)/2$ with $c_{s}=0$. As the mesons propagate in both directions, the center-of-mass coordinate’s distribution widens around the initial state average. The width is characterized by $c_{s}$ and is expected to increase as a function of time. Its slope gives the speed of propagation of excitations.
In Fig. 3(a),(b), we show $r_{\text{avg}}$ and $c_{s}$ respectively as a function of time for a typical case when $h=1.1J$. After some transient behavior, $r_{\text{avg}}$ settles and oscillates around a steady value of $r^{\prime}_{\text{avg}}=1.46$. This suggests that the propagating excitations are a superposition of different meson sizes. We also see that $r_{\text{avg}}$ oscillates in time, suggesting meson size oscillations. The width of center-of-mass distribution $c_{s}$ steadily increases from $c_{s}=0$ as the mesons propagate outwards from the center. This is expected since the confined mesons can freely delocalize. The average size of propagating mesons and frequency of size oscillations are strongly dependent on the confining field, $h/J$.
Figure 3: Dynamics after initializing an $r=1$ meson at the center of a chain of $L=100$ sites for $h=1.1J$. (a) Average size of propagating mesons, $r_{\text{avg}}$, vs time, $t$. The meson size oscillates in time about a value of $r^{\prime}_{\text{avg}}=1.46$. (b) Width, $c_{s}$, of the center-of-mass distribution vs time, $t$. Due to propagating mesons, the width of the center-of-mass distribution increases monotonically with time.
In Fig. 4, we plot $r^{\prime}_{\text{avg}}$ as a function of the confinement field $h/J$. This is the value around which $r_{\text{avg}}$ oscillates in time. We find that as $h/J$ decreases, the propagating meson excitations get longer such that $r^{\prime}_{\text{avg}}-1\propto 1/h$. We also extract the frequency of oscillations of $r_{\text{avg}}$ at long time. We take the Fourier transform of the $r_{\text{avg}}$ vs time data at long times and use the frequency of the dominant peak as the oscillation frequency. Fig. 4 shows this frequency of oscillations, $\omega$ as a function of $h/J$. We find that $\omega\propto h$ for large $h/J$ and deviates for smaller $h/J$.

### v2

@@ Abstract
Quantum simulation of lattice gauge theories is an important avenue to gain insights into both particle physics phenomena and constrained quantum many-body dynamics. There is a growing interest in probing analogs of high energy collision phenomena in lattice gauge theories that can be implemented on current quantum simulators. Motivated by this, we characterize the confined mesons that originate from a local high energy excitation in a particle-conserving 1D $Z_{2}$ lattice gauge theory. We focus on a simple, experimentally accessible setting that does not require preparation of colliding wavepackets and isolates the effects of gauge field confinement strength and initial state energy on the nature of propagating excitations. We find that the dynamics are characterized by the propagation of a superposition of differently sized mesons. The linear confinement leads to meson size oscillations in time. The average meson size and oscillation frequency are controlled by the strength of the gauge field confinement. At a constant confinement field, the average meson length is controlled by the initial excitation’s energy. Higher energies produce longer mesons and their effective mass depends strongly on their size: longer mesons propagate more slowly out of the central excitation. Mesons of different sizes get spatially filtered with time due to different speeds. We show that this phenomenology is a consequence of linear confinement and remains valid in both the strong and weak confinement limit. We present simple explanations of these phenomena supported by exact numerics.
@@ I Introduction
Quantum computers and quantum simulators can bridge this gap. By implementing these gauge theories on a discrete lattice, error corrected quantum computers would be able to simulate real-time quantum dynamics of high-energy phenomena to high levels of precision. Current quantum processors can already implement dynamical simulation experiments of some simple models such as $Z_{2}$ and $U(1)$ lattice gauge theories, albeit with significant noise effects [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
@@
The study of dynamical phenomena in lattice gauge theories is attracting considerable attention in recent years. These gauge theories show exotic non-ergodic phenomena such as lack of thermalization due to confinement and Hilbert-space fragmentation [31, 32, 33, 34, 8, 35], quantum many-body scars [36, 37, 38, 39] and Floquet time crystals [40]. Analogs of QCD phenomena such as confinement dynamics [41, 14, 13, 42], pair production through string breaking [43, 44, 17, 45, 46, 47, 15, 16, 12], false vacuum decay [48], and dynamical hadron formation [49, 50] have been found in theoretical calculations and some quantum simulator experiments.
These successes have opened up an exciting new avenue of probing analogs of high energy particle collision phenomena within these simple lattice gauge theories. Some recent works have studied schemes to prepare moving wavepackets for collisions  [51, 52, 53] and numerically analyzed scattering processes [53, 54, 55, 49]. In these studies, the dynamics involves a rich interplay of multiple kinds of bound excitations, resonant processes of particle pair production and string fluctuation dynamics.
@@
We also consider the effect of tuning the initial state energy at a fixed confinement field. The average length of the meson that propagates monotonically increases as a function of initial state energy, instead of the spatial size of the initial excitation. As the meson size increases, its propagation speed is slower, which slows down the spread of particle density away from the excitation region. Since the long time wavefunction of even a single meson is a superposition of differently sized mesons, the correlation between meson size and speed causes a size dependent spatial separation of the mesons participating in the superposition. This is a consequence of linear confinement and we find that this feature holds true even for cases where the confining field is weaker than the particle hopping term. We show that for weaker confining fields, mesons longer than a certain size experience a slow down.
@@ II Model
Figure 1: (a) Meson dynamics: A meson propagates out of the localized energy excitation in both directions in a 1D chain governed by a $Z_{2}$ lattice gauge theory. The state is described by a superposition of different meson sizes. The confinement field and the initial excitation energy determine the meson size distribution. The propagation speed varies with the meson size. (b) A basis state of the 1D $Z_{2}$ lattice gauge theory model. Open and filled circles represent absence and presence of a boson on sites labeled by index ii. The red and blue arrows are link gauge spins representing $\sigma^{z}=\pm 1$ respectively. The quantity, $G_{i}=\sigma^{z}_{i-1,i}(-1)^{n_{i}}\sigma^{z}_{i,i+1}=1$ is conserved on each site. The red ellipse denotes a meson where the two bosons are next to each other, called an $r=1$ meson.
@@
There is a local conservation law at each site since $G_{i},H]=0$. The Hamiltonian eigenstates can thus be chosen to also be the eigenstates of the operator $G_{i}$ at each site. The eigenvalues of $G_{i}$ are $\pm 1$. Different choices of the $G_{i}$ eigenvalues correspond to independent and physically distinct sectors of the Hamiltonian’s Hilbert space [24]. Throughout this work, we only consider states $|\psi\rangle$ where $G_{i}|\psi\rangle=+1|\psi\rangle$ for all sites $i$. This constraint implies that $\sigma^{z}_{i-1,i}(-1)^{n_{i}}\sigma^{z}_{i,i+1}|\psi\rangle=|\psi\rangle$. We can rewrite it as, $\sigma^{z}_{i-1,i}\sigma^{z}_{i,i+1}|\psi\rangle=e^{i\pi/2(\sigma^{z}_{i-1,i}-\sigma^{z}_{i,i+1})}=(-1)^{n_{i}}|\psi\rangle$. This can be interpreted as a Gauss law where the local matter density determines the electric field configuration around it.
@@ III The two-particle sector
The Hamiltonian in Eq. 1 conserves total boson number. In this work, we consider dynamics when there are only two bosons present in the system. We consider configurations where the two bosons are added onto the vacuum, the particle-free ground state. The vacuum state has all the gauge spins aligned along the negative $z$-axis ($\sigma^{z}_{i,i+1}=-1$) to minimize the electric field energy (we assume $h>0$). An example of such a two-particle configuration is shown in Fig. 1(b). Due to the Gauss law constraint on each site, the link spin configuration can be completely determined by the position of the bosons. We thus get, $\sigma^{z}_{i,i+1}=\prod_{j\leq i}(-1)^{n_{j}}$. The states in the two-boson sector can be characterized by the position indices of the two bosons, $i_{1}$ and $i_{2}$. The link spins between sites $i_{1}$ and $i_{2}$ are the only ones aligned in the positive $z$-direction leading to an electric field energy cost of +2h+2h for each such link. This is what leads to confinement of particles into mesons in this model.
@@
$H=-J\sum_{i_{1}\neq i_{2}}(|i_{1}+1,i_{2}\rangle\langle i_{1},i_{2}|+|i_{1},i_{2}+1\rangle\langle i_{1},i_{2}|+\text{h.c.})\\ +2h\sum_{i_{1}\neq i_{2}}|i_{1}-i_{2}||i_{1},i_{2}\rangle\langle i_{1},i_{2}|.$ (2)
Due to the hard-core limit for the bosons, the position coordinates for the two bosons must be unequal. We use the center-of-mass $(c=(i_{1}+i_{2})/2)$ and relative $(r=i_{1}-i_{2})$ coordinates to further simplify the Hamiltonian. The Hamiltonian in Eq. 2 becomes
$\begin{split}H&=-J\sum_{c}\sum_{r=1}^{\infty}(|c+1/2\rangle\langle c|+\text{h.c.})(|r+1\rangle\langle r|+\text{h.c.})\\ &+2h\sum_{r=1}^{\infty}r|r\rangle\langle r|\end{split}$ (3)
Here the center of mass coordinate $c$ can take half-integer and integer values while the relative coordinate $r$ only takes integer values where $r>1$ due to the hardcore boson limit. The relative coordinate has a linear potential term in Eq. 3. In the ground state, at any finite $h/J$, two particles are confined by this potential into mesons [56]. The center-of-mass coordinate is free and thus the mesons can delocalize. In this work, we study the two-particle sector dynamics in terms of the center-of-mass and relative coordinates. The different relative coordinate values correspond to different types of possible meson excitations. The smallest possible meson is the $r=1$ meson where the two bosons are right next to each other. In later sections, we calculate how the size of the propagating mesons depends on field term $h/J$ and the energy of the initial local excitation. We will also determine how meson size affects its propagation speed.
@@ III.1 Diagonalizing the Hamiltonian
We can diagonalize the Hamiltonian in Eq. 3 by considering the eigenstates to be of the form, $|\psi_{n,k}\rangle=\Phi_{k}\otimes\Gamma_{n}=(\sum_{c}\phi^{k}_{c}|c\rangle)\otimes(\sum_{r}\gamma^{n}_{r}|r\rangle)$ that satisfy the equation, $H|\psi_{n,k}\rangle=E_{n,k}|\psi_{n,k}\rangle$. Note that the center-of-mass states only have a hopping term in the Hamiltonian. Thus plane wave states of the form, $\phi^{k}_{c}=e^{ikc}/\sqrt{2L-1}$ are eigenstates for the center-of-mass degree of freedom. Here $L$ is the total number of sites in the chain. Due to the half integer steps of the coordinate $c$, the momentum $k$ ranges from $2\pi$ to $2\pi$ in steps of $2\pi/(2L-1)$. We can find the function $\gamma^{n}_{r}$ by substituting $\phi^{k}_{c}$ into the eigenvalue equation of the Hamiltonian in Eq. 3 and projecting it onto $|c\rangle\otimes|r\rangle$. We then get
@@
Note that the relative coordinate is restricted to $r\geq 1$. Without such a restriction, this describes a Wannier-Stark ladder and Eq. 4 admits Bessel function solutions where $\gamma^{n}_{r}=J_{r-n}(2J\cos(k/2)/h)$ [57]. The energy eigenvalue in that case is given by, $E_{n,k}=2hn$ where $n$ is integer quantized. The energy is independent of the center-of-mass momentum leading to a large degeneracy. However in our case, $r\leq 0$ is not allowed and thus we must enforce the boundary condition that $\gamma^{n}_{r=0}=0$. This is akin to adding an infinite potential wall at $r=0$. For a semi-infinite chain with a hard wall at $r=0$, the eigenstate solutions are again Bessel functions where $n$ is no longer integer quantized [35, 53]. The value of $n$ is found from the condition that at $r=0$, we must have $J_{-n}(2J\cos(k/2)/h)=0$. For finite values of $h/J$, we would find a non-integer momentum dependent value of $n$ such that $2J\cos(k/2)/h$ is a zero of the Bessel function. The energy bands would thus no longer be flat with respect to momentum $k$. When $h/J\to\infty$, $n$ tends towards integer values independent of $k$ and we recover the flat Wannier-Stark bands. If the chain is finite with hard wall boundaries at both ends, the exact eigenstates are given by Lommel polynomials and the eigenenergies are obtained from the zeros of the Lommel polynomial [58].
Even with these boundary conditions, the eigenstates and eigenenergies are analytically rather simple in the large-$h/J$ and small-$h/J$ limit. These limits essentially capture a complete picture of the physics. We now derive the solution in these two simple limits.
@@ III.1.1 Large-$h/J$ limit
When $h/J\gg 1$, the Bessel function solutions with integer quantized eigenenergies approach the exact eigenstates. At large $h/J$, the asymptotic form of the solutions is, $J_{r-n}\sim(2J\cos(k/2)/h)^{|r-n|}$. Since $h/J\gg 1$, the amplitude shows a decay with an increasing value of the exponent, |rn||r-n|. The amplitude at $r=0$ is $\sim(2J\cos(k/2)/h)^{|n|}$. As $h/J\to\infty$, when $n\geq 1$, this amplitude goes to zero, satisfying the boundary condition in the asymptotic limit. The energy eigenvalues are quantized in this limit and given by, $E_{n,k}=2hn$. The $n^{\text{th}}$ eigenstate is sharply peaked around $r=n$ with power law tails ($1/h^{|r-n|}$). Seen another way, the zeroth order $h/J\to\infty$ solution has the $n^{\text{th}}$ eigenstate as a delta function on $r=n$ with the eigenenergy $2hn$ where $n\geq 1$. The hopping term proportional to $J$ in Eq. 4 provides first order perturbation to the wavefunction and gives it a tail such that the amplitude on level $r\neq n$ is $(2J\cos(k/2)/h)^{|r-n|}$.
@@ III.1.2 Small-$h/J$ limit
$\frac{d^{2}\gamma^{n}_{r}}{dr^{\prime 2}}+(r^{\prime}-E^{\prime}_{n,k})\gamma^{n}_{r}=0$ (6)
@@
At intermediate values of the confining field $h/J$, one needs to find the exact momentum dependent eigenenergies using the zeros of the non-integer order Bessel functions [35, 53].
@@ IV Effect of confinement field on meson dynamics
In this section, we characterize the meson dynamics from simple initial states as a function of confinement field strength. We first consider the initial excitation where an $r=1$ meson is localized at the center of a 1D chain of length $L$. This corresponds to the initial configuration shown in Fig. 1(b). In the $|r,c\rangle$ (relative and center-of-mass coordinate) basis, the wavefunction at time $t=0$ is given by $|\psi(0)\rangle=|1,(L+1)/2\rangle$. Under the Hamiltonian in Eq. 3, the system wavefunction evolves in time and can be written as, $|\psi(t)\rangle=\sum_{c,r}g_{c,r}(t)|c,r\rangle$. In our calculations, we fix system size to $L=100,120$ and use exact diagonalization to calculate the time-evolved wavefunctions. Appendix D discusses details of this calculation.
@@
Figure 2: (i),(ii): Particle density, $n$ as a function of site index (x-axis) and time, JtJt (y-axis) for (i) $h/J=1.1$ and (ii) $h/J=0.3$. With time, the meson delocalizes in light-cones like structures. The outermost cones represent faster moving excitations. The slower moving inner cones show substructures with diamond shaped particle tracks denoting temporal oscillation of meson size. (iii),(iv): Snapshot of occupation probability, $p$ as a function of meson size ($r$) and its center-of-mass position ($c$) at (iii) $Jt=50$ for $h=1.1J$ and (iv) $Jt=30$ for $h=0.3J$. The red dashed lines in (i),(ii) denotes these two chosen time steps. In both cases, smaller sized mesons delocalize the farthest, suggesting larger mesons propagate slower. The different speeds lead to a spatial filtering of meson sizes as such that it is more probable to detect a longer meson closer to the center during time evolution. Note that we have removed zero density pixels to correct for a parity effect in (iii),(iv) because when $r$ is even, the center-of-mass $c$ only takes integer values while for odd $r$, it only takes half-integer values.
@@
We show the dynamics of propagating excitations in Fig. 2 at two values of the confining field, $h=1.1J$ and $h=0.3J$. In Fig. 2(i),(ii), we plot the expectation value of the particle number, $n$, as a function of site index and time, $Jt$. The particles delocalize through light-cone like structures emanating from the central excitation region at time $Jt=0$. The outermost cones in both cases represent fast moving excitations. The inner cone has a slower spread and has substructures that show diamond shaped particle tracks that also slowly spread out from the center. These are more pronounced and wider for smaller $h=0.3J$. These tracks show similarities to the ones found after wavepacket collisions in Ref. [55, 51] where they were interpreted as the two particles coming together and colliding multiple times. We will show in later sections that we can interpret them as meson size oscillations which gradually spread out from the center as these mesons delocalize. Similar oscillations were also observed in studies involving Hamiltonian quench dynamics in Ising models [59, 13, 41]. In Fig. 2(iii),(iv), we plot the occupation probability, $p$, of a meson of length $r$ occupying a center-of-mass position $c$. We can clearly see that at a given time, the largest center-of-mass positions away from the central excitation region have the highest probability of occupation from a smaller sized meson. This suggests that meson size is inversely related to propagation speed. As the long time wavefunction of a single meson is characterized by a superposition of different meson sizes, this speed differential causes spatial separation of the mesons involved in the superposition according to their size. We will show that this is a consequence of the linear confinement of mesons and persists for both $h>J$ and $h<J$ regimes.
@@
We use the system wavefunction at long times to characterize the propagating excitation. In particular, we calculate the average meson size, $r_{\text{avg}}(t)=\sum_{r}r|\langle r|\psi(t)\rangle|^{2}$ as a function of time. The meson excitation propagates equally in both the directions away from the center. We determine the speed at which the excitation propagates by calculating the temporal slope of the center-of-mass coordinate’s standard deviation, $c_{s}$, defined as
@@
At time $t=0$, the average center-of-mass coordinate is at $(L+1)/2$ with $c_{s}=0$. As the excitation propagates in both directions, the center-of-mass coordinate’s distribution widens around the initial state average. The width is characterized by $c_{s}$ and is expected to increase as a function of time. Its slope gives the speed of propagation.
In Fig. 3, we show $r_{\text{avg}}$ and $c_{s}$ as a function of time for two values of $h/J$, $h=1.1J$ (Fig. 3(a,b)) and $h/J=0.3$ Fig. 3(c,d). After some transient behavior, $r_{\text{avg}}$ settles and oscillates around a steady value of $r^{\prime}_{\text{avg}}=1.46$ and $r^{\prime}_{\text{avg}}=3.28$ for $h=1.1J$ and $h=0.3J$ respectively. This suggests that the propagating excitation is a superposition of different meson sizes. We also see that $r_{\text{avg}}$ oscillates in time, suggesting meson size oscillations. The oscillation amplitude is expected to asymptote to a constant value in the infinite time limit. A slight decay of this amplitude even at longer times in our plots is a consequence of our finite time simulations. The finite system size in our simulation restricts us from going to longer times as the particles start hitting the boundary.
The width of center-of-mass distribution $c_{s}$ steadily increases from $c_{s}=0$ as the excitation propagates outwards from the center. This is expected since confined mesons can freely delocalize. This delocalization is faster for $h=0.3J$ compared to $h=1.1J$. The average size of the propagating meson and frequency of size oscillations are strongly dependent on the confining field, $h/J$.
Figure 3: Dynamics after initializing an $r=1$ meson at the center of a chain of $L=100$ and $L=120$ sites for $h=1.1J$ and $h=0.3J$ respectively. (a,c) Average size of propagating mesons, $r_{\text{avg}}$, vs time, $t$. The meson size oscillates in time about a value of $r^{\prime}_{\text{avg}}=1.46$ and $r^{\prime}_{\text{avg}}=3.28$ for $h/J=1.1J$ and $h=0.3J$ respectively. (b,d) Width, $c_{s}$, of the center-of-mass distribution vs time, $t$. Due to the propagating meson, the width of the center-of-mass distribution increases monotonically with time. The width grows faster for lower value of $h/J$.
In Fig. 4, we plot $r^{\prime}_{\text{avg}}$ as a function of the confinement field $h/J$. This is the value around which $r_{\text{avg}}$ oscillates in time. We find that as $h/J$ decreases, the propagating meson excitation gets longer such that $r^{\prime}_{\text{avg}}-1\propto J/h$. The error bars show the long time average of the standard deviation of meson size, $r_{\text{sd}}^{\prime}$. We first find the standard deviation of meson size, $r_{\text{sd}}(t)=\sum_{r}\sqrt{r^{2}|\langle r|\psi(t)|^{2}-r_{\text{avg}}(t)^{2}}$. The long time average of $r_{\text{sd}}(t)$ gives $r_{\text{sd}}^{\prime}$. As $h/J$ decreases, this quantity increases as seen in the larger error bars in Fig. 4, suggesting larger fluctuations in meson size.
We also extract the frequency of oscillations of $r_{\text{avg}}$ at long time. We take the Fourier transform of the $r_{\text{avg}}$ vs time data at long times and use the frequency of the dominant peak as the oscillation frequency. Fig. 4 shows this frequency of oscillations, $\omega$ as a function of $h/J$. We find that $\omega\propto h$ for large $h/J$ and deviates for smaller $h/J$. Note that as $h/J$ reduces, the finite size effects become more important as the meson undergoes large fluctuations in size with faster delocalization towards the boundary of our system.